OpenStudy (anonymous):

Kenji has 2500 Japanese yen. If 1 US dollar equals 90.39 Japanese yen, about how many dollars can Kenji buy for his yen?

OpenStudy (whpalmer4):

\[2500 \text{ yen} * \frac{1 \text{ dollar}}{90.39 \text{ yen }} = \text{???}\text{ dollar}\]

OpenStudy (anonymous):

Kenji can buy $27.65 with his 2500 yen but then 2500/90.39=27.66

OpenStudy (calculusxy):

snce you want to know how much usd is needed, you want to dIvIde the amount ofyen by the amount of usdfor one yen

OpenStudy (anonymous):

Exactly ^

OpenStudy (whpalmer4):

@calculusxy @angelicmelody no, that calculation is incorrect. you'd be dividing \[\frac{\text{yen}}{\frac{\text{usd}}{\text{yen}}}\] which gives you \[\frac{\text{yen}^2}{\text{usd}}\] Dimensional analysis is your friend.

OpenStudy (anonymous):

ah yeah i did do it wrong.. gomenasai im half japanese myself

OpenStudy (anonymous):

gomenasai means im very sorry/sorry ect btw @whpalmer4

OpenStudy (whpalmer4):

the first one you did is right — it's equivalent to mine. My sweetie's family name is Murakami, so I've encountered "gomenasai" before :-)